OpenStudy (shubhamsrg):

Please have a look at this geometry question. In the mean time, I am posting my attempt.

OpenStudy (shubhamsrg):

http://gyazo.com/cb97188aea63e902b7547862885ac945

OpenStudy (shubhamsrg):



OpenStudy (shubhamsrg):

a= 1/2^n b=1/2^(n+1) we have to solve for x

OpenStudy (shubhamsrg):



OpenStudy (shubhamsrg):

OA = a+x OB=b+x AB=a+b DO= sqrt((a+x)^2 - (a-x)^2) = 2sqrt(ax) and same way OC = 2sqrt(bx)) Now, AB^2 = OA^2 + OB^2 - 2OA.OB cos(<BOA>) => (a+b)^2 = (a+x)^2 + (b+x)^2 + 2(a+x)(b+x)cos<DOA + COB> => (a+b)^2 = (a+x)^2 + (b+x)^2 + 2(a+x)(b+x) [ cosDOA cosCOB - sinDOA sinCOB ] => (a+b)^2 = (a+x)^2 + (b+x)^2 + 2(a+x)(b+x) [ 2sqrt(ax)/(a+x) . 2sqrt(bx)/(b+x) - (a-x).(b-x)/(a+x)(b+x) ] Is everything right so far?

OpenStudy (shubhamsrg):

oh, seems like I got my mistake! :O